The Si- ion is stable with respect to the neutral atom and an electron. This is evident in the value of the electron affinity of silicon which is 133.6 kJ mol-1. What is the longest wavelength of light that could remove an electron from a Si- ion? In which region of the electromagnetic spectrum does this light fall?

Wavelength = ? nm

Region = ____?_____ultra violet/visible/infrared
Answer #1

energy of one mole photon = NA * h c / wavelength

or

133.6 * 1000 = 6.023 * 10^23 * 6.626 * 10^-34 * 3.0 * 10^8 / wavelength

or

wavelength = 8.96 * 10^-7 m = 896 * 10^-9 m = 896 nm

thus

longest wavelength of light = 896 nm

896 nm fall in 800-1400 nm range.

Hence

Region = infrared